OpenStudy (anonymous):

Solve bx – cx = 7 for x.

OpenStudy (anonymous):

1st factor out an x on the left side.

OpenStudy (anonymous):

A. X=7(b-c) B. X=7/b - c C. X=b-c + 7 D. X=7/b-c

OpenStudy (anonymous):